OpenStudy (anonymous):

The formula P=0.68x^2 -0.041x+3 models the approx. population in thousands, for a species of fish in a local pond, x years after 1997.During what year will the population reach 27,234? I know how to set up the equation i just keep getting the wrong answer...the answer is 2003, but im not sure how to get to that

OpenStudy (helder_edwin):

u r supposed to solve \[ \large 27234=0.68x^2-0.041x+3 \] right??

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

rearranging u get \[ \large 0.68x^2-0.041x-27231=0 \] if u use the quadratic formula u have \[ \large x=\frac{0.041\pm\sqrt{(-0.041)^2-4\cdot0.68(-27231)}}{2\cdot0.68} \] right??

OpenStudy (anonymous):

yeah

OpenStudy (helder_edwin):

u got \[ \large x_1\approx-200.084 \] \[ \large x_2\approx200.144 \] right??

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

x is the number of years after 1997 so the answer would be \[ \large 1997+x_2=1997+200=2197 \]

OpenStudy (helder_edwin):

there must be something wrong with the wording of the problem then

OpenStudy (anonymous):

Yeah Ive been trying to do this problems for forever so I totally agree lol. Thanks though!

OpenStudy (helder_edwin):

u r welcome